Use work and energy to find an expression for the
Answer #1

1. The block M moves distance h
Work done by friction = mu*M*g
Energy lost by the m = m*g*h
m*g*h - mu*M*g*h = 0.5(M+m)v^2
v = sqroot(2*gh(m - mu*M)/(M+m))
2. For frictionless table, mgh = 0.5(M+m)v^2
V = sqroot(2mgh/(M+m))
Answer #2

dumb dumb dumb dumb dumb dumb dumb dumb dumb dum dbum dumdb dumb dumb dumb dumb dumb dumb udmb umdb udmb udmb dum bdumb dumb dumb dumb udmb

source: mama
answered by: joe